OpenStudy (anonymous):

Reference number for 5/8 pi

OpenStudy (dumbcow):

reference angles are between 0 and pi/2 5pi/8 is in quadrant 2 so ref angle = pi - 5pi/8 = 3pi/8

OpenStudy (anonymous):

\[\frac{ 5\pi }{ 8 }?\]

OpenStudy (dumbcow):

link to unit circle study it , use it as a reference for these type of questions http://en.wikipedia.org/wiki/File:Unit_circle_angles_color.svg